OpenStudy (anonymous):

Medal for help please...simplify (a^2 + 2a - 15) / (a^2 - 16) / (9 - 3a) / (3a - 12)

OpenStudy (anonymous):

@iheartmath123 : Sorry but Can't get your qustn please Re write using Equation or drawing tab..

OpenStudy (anonymous):

\[\frac{a^2 + 2a - 15}{ a^2 - 16} \div \frac{ 9 - 3a }{3a - 12}\]

OpenStudy (anonymous):

this is the answer i got but not sure if it is correct... \[-\frac{ a + 5 }{ a - 4? }\]

OpenStudy (anonymous):

you have done little mistake.. i thik so.. in denomenator it will come (a+4)

OpenStudy (anonymous):

oh yes, thank you i typed it in incorrectly...should be \[-\frac{ a + 5 }{a + 4}\]...is that what you got?

OpenStudy (anonymous):

yuup .. wll done.. you dont have any problem now ... :)

OpenStudy (anonymous):

thanks so much, i really appreciate it!

OpenStudy (anonymous):

wlcome... :)

OpenStudy (anonymous):

close the qstn once you got it.